# Large Language Models as Neurolinguistic Subjects: Discrepancy in Performance and Competence for Form and Meaning

### v2

@@ 7 Conclusion
Appendix A Extended Cognitive Science Background
@@
Mesgarani and Chang (2012) is a representative work employing decoding analysis in the realm of language and speech. They use electrocorticography (ECoG) to record neural responses from subjects who are listening to speech. Leveraging decoding analysis, they were able to differentiate between neural patterns induced by attended speech and those elicited by background speech, (to be ignored by the test persons), thereby highlighting the perceptual differences between the two speech stimuli conditions.
@@ B.1 Form: BLiMP, CLiMP, and DistilLingEval
CLiMP (Xiang et al., 2021) is a corpus of Chinese grammatical minimal pairs consisting of 16 datasets, each containing 1,000 sentence pairs. CLiMP covers 9 major Chinese language phenomena in total, less than the BLiMP dataset due to the less inflectional nature of Mandarin Chinese. The vocabulary of the CLiMP dataset is based on the translation of the BLiMP dataset, with words and features specific to Chinese added.
@@ B.2 Meaning: COMPS, COMPS-ZH, and COMPS-DE
In summary, out of 521 concepts, manual corrections were made to 57 entries in the Chinese dataset and 49 in the German dataset. Similarly, from 3,592 properties, 713 required manual corrections in the Chinese dataset, and 512 in the German dataset. This rigorous process was essential for preserving the integrity and reliability of the translated datasets.
@@ Appendix C Model Details
In our experiments, we used three open-source LLMs, two English-centric LLMs (Llama2 and Llama3), and one multilingual LLMs (Qwen) with a focus on English and Chinese. These models were trained on different amounts of English, Chinese, and German data (see Table 3).
@@
Llama2 (Touvron et al., 2023b) and Llama3 (AI, 2024) are two English-centric LLMs which represent an advanced iteration of the Llama foundation models developed by Meta AI (Touvron et al., 2023a). The Llama madels were trained on publicly available corpora predominantly in English. Despite this focus, Llama models are also exposed to a limited amount of multilingual data. Llama 1, for example, is pretrained on an extensive scale of corpora comprising over 1.4 trillion tokens, of which less than 4.5% constitute multilingual data from 20 different languages. Llama 2 expands this linguistic diversity, featuring 27 languages each representing more than 0.005% of the pertaining data. Therefore, English-centric models harness multilingual abilities (Lai et al., 2023). In this work, we use Llama2-7B and Llama3-8B for our experiments.
@@
QWen is a series of LLMs developed by Alibaba Inc. Bai et al. (2023). Qwen was trained on 2-3 trillion tokens of multilingual pretaining data. It is essentially a multilingual LLM with the focus on English and Chinese. We use the Qwen-7B model in our experiments.
Appendix D Extended Neurolinguistic Results for English
@@
1) For the ellipsis task especially, local features in layer 0 (the embedding layer before the Transformer structure) already provide sufficient linguistic information to accomplish the task without any contextual information.
@@
Compared to Llama2, Llama3 won’t improve internal grammatical capabilities much but will learn concepts better and faster.
@@
Figure 7: T-test results between Llama2 and Llama3 feature learning saturation layer. The symbols ‘***’, and ‘*’ denote t-test p-values less than 0.001, and 0.05, respectively.
Figure 7 compares the feature learning saturation layers between Llama2 and Llama3. The results for form learning (blue bars) do not differ significantly between Llama2 and Llama3 and weakly significantly between Llama2_chat and Llama3_chat. However, the results for meaning learning (blue bars) are both highly significant, indicating that Llama3 requires fewer layers to encode conceptual features than Llama2. This suggests that Llama3 comprehends sentences faster.
Appendix E Extended Neurolinguistic Results for Multilingual Analysis
1) Llama’s performance on Chinese. Despite Chinese not being the primary training language of the Llama2 models, they still perform well in encoding Chinese form/grammar. However, both Qwen, which is primarily trained on Chinese, and the Llama models show relatively poor performance in understanding Chinese meaning/concepts.
2) Improvement in Llama3 for Chinese Semantics. Llama3 shows some improvement over Llama2 in understanding Chinese semantics, as indicated by the slightly higher red curve in the middle row. The improvement in syntactic understanding is minimal.
3) Qwen’s Faster Syntax Learning but Slower Semantic Encoding for Chinese. Compared to the Llama models, Qwen learns Chinese grammar faster, as indicated by the sharper rise of the blue curve. However, it encodes Chinese semantics more slowly, evidenced by the larger gap between the form and meaning curves in the early layers.
4) Poor Performance for German. For German, a low-resource language, all three models perform poorly. Despite Chinese not being a primary training language for the Llama models, their performance is relatively decent, suggesting that the actual proportion of German training data might be much smaller. This highlights differences in the resource allocation for the three languages. 5) Form needs less data to capture compared to meaning. From Table 3, Chinese is classified as a mid-resourced language for Llama, yet it achieves high form competence (but low meaning competence), suggesting that capturing form requires less data than meaning.
Appendix F Extended Discussion

### v3

@@ 7 Conclusion
Acknowledgements
We thank the anonymous reviewers for their valuable advice and feedback. This research was supported by DFG (German Research Foundation) grant SCHU 2246/14-1 and Munich Center for Machine Learning (MCML).
Appendix A Neuro Probing’s Cognitive Science Background
@@
Mesgarani and Chang (2012) is a representative work employing decoding analysis in the realm of language and speech. They use electrocorticography (ECoG) to record neural responses from subjects who are listening to speech. Leveraging decoding analysis, they were able to differentiate between neural patterns induced by attended speech and those elicited by background speech (to be ignored by the test persons), thereby highlighting the perceptual differences between the two speech stimulus conditions.
@@ B.1 Form: BLiMP, CLiMP, and DistilLingEval
CLiMP (Xiang et al., 2021) is a corpus of Chinese grammatical minimal pairs consisting of 16 datasets, each containing 1,000 sentence pairs. CLiMP covers 9 major Chinese language phenomena in total, fewer than the BLiMP dataset due to the less inflectional nature of Mandarin Chinese. The vocabulary of the CLiMP dataset is based on the translation of the BLiMP dataset, with words and features specific to Chinese added.
@@ B.2 Meaning: COMPS, COMPS-ZH, and COMPS-DE
In summary, out of 521 concepts, manual corrections were made to 57 entries in the Chinese dataset and 49 in the German dataset. Similarly, out of 3,592 properties, 713 required manual corrections in the Chinese dataset, and 512 in the German dataset. This rigorous process was essential for preserving the integrity and reliability of the translated datasets.
@@ Appendix C Model Details
In our experiments, we used three open-source LLMs, two English-centric LLMs (Llama2 and Llama3), and one multilingual LLM (Qwen) with a focus on English and Chinese. These models were trained on different amounts of English, Chinese, and German data (see Table 3).
@@
Llama2 (Touvron et al., 2023b) and Llama3 (AI, 2024) are two English-centric LLMs which represent an advanced iteration of the Llama foundation models developed by Meta AI (Touvron et al., 2023a). The Llama madels were trained on publicly available corpora predominantly in English. Despite this focus, Llama models are also exposed to a limited amount of multilingual data. Llama 1, for example, is pretrained on an extensive scale of corpora comprising over 1.4 trillion tokens, of which less than 4.5% constitute multilingual data from 20 different languages. Llama 2 expands this linguistic diversity, featuring 27 languages, each representing more than 0.005% of the pertaining data. Therefore, English-centric models harness multilingual abilities (Lai et al., 2023). In this work, we use Llama2-7B and Llama3-8B for our experiments.
@@
QWen is a series of LLMs developed by Alibaba Inc. Bai et al. (2023). Qwen was trained on 2-3 trillion tokens of multilingual pre-training data. It is essentially a multilingual LLM with a focus on English and Chinese. We use the Qwen-7B model in our experiments.
Appendix D Supplemented Results for English
@@
1) For the ellipsis task, especially, local features in layer 0 (the embedding layer before the Transformer structure) already provide sufficient linguistic information to accomplish the task without any contextual information.
@@
Compared to Llama2, Llama3 won’t improve internal grammatical capabilities much, but will learn concepts better and faster.
@@
Figure 13: T-test results between Llama2 and Llama3 feature learning saturation layer. The symbols ‘***’, and ‘*’ denote t-test p-values less than 0.001 and 0.05, respectively.
Figure 13 compares the feature learning saturation layers between Llama2 and Llama3. The results for form learning (blue bars) do not differ significantly between Llama2 and Llama3, and weakly significantly between Llama2_chat and Llama3_chat. However, the results for meaning learning (blue bars) are both highly significant, indicating that Llama3 requires fewer layers to encode conceptual features than Llama2. This suggests that Llama3 comprehends sentences faster.
Appendix E Supplemented Results for Multilingual Analysis
Figure 14: Correlation between meaning competence and form competence.
Meaning competence is correlated to form competence.
Figure 14 illustrates the relationship between Form competence (x-axis) and Meaning competence (y-axis) across English, German, and Chinese in the neuro assessment for LLMs. The positive correlation (R² = 0.48) suggests that higher meaning competence generally corresponds to higher form competence.
Llama’s performance on Chinese.
Despite Chinese not being the primary training language of the Llama2 models, they still perform well in encoding Chinese form/grammar. However, both Qwen, which is primarily trained on Chinese, and the Llama models show relatively poor performance in understanding Chinese meaning/concepts.
Improvement in Llama3 for Chinese Semantics.
Llama3 shows some improvement over Llama2 in understanding Chinese semantics, as indicated by the slightly higher red curve in the middle row. The improvement in syntactic understanding is minimal.
Qwen’s Faster Syntax Learning but Slower Semantic Encoding for Chinese.
Compared to the Llama models, Qwen learns Chinese grammar faster, as indicated by the sharper rise of the blue curve. However, it encodes Chinese semantics more slowly, evidenced by the larger gap between the form and meaning curves in the early layers.
Poor Performance for German.
For German, a low-resource language, all three models perform poorly. Despite Chinese not being a primary training language for the Llama models, their performance is relatively decent, suggesting that the actual proportion of German training data might be much smaller. This highlights differences in the resource allocation for the three languages.
Form needs less data to capture compared to meaning.
From Table 3, Chinese is classified as a mid-resourced language for Llama, yet it achieves high form competence (but low meaning competence), suggesting that capturing form requires less data than meaning.
Appendix F Supplemented Discussion